Jack weighs 4/5 of his weight plus 40 lbs. Write the algebraic equation and solve. OK, so I know the answer, just by looking at the problem, but my teacher requires that I write the equation and explain, which I cannot do. Please if anybody can help me understand how to do this, I would appreciate it. I struggle with how to write the proper equations.
Answer by scott8148(6628)

the semantics is usually the problem ; instead of saying "Jack weighs" , say "Jack's weight is"

Jack's weight ___ W

is ___ =

4/5 of his weight plus 40 lbs. ___ (4/5)W + 40

W = (4/5)W + 40

subtracting (4/5)W ___ (1/5)W = 40

multiplying by 5 ___ W = 200